National Defence Academy (NDA) · Topic Collection

Quantitative Aptitude PYQs

Arithmetic, number systems, percentages, ratios, profit and loss, time and work, algebra, geometry, mensuration, and trigonometry.

2520 Total Questions21 Official PapersYears: 2015–2025Search in Quantitative Aptitude
Showing questions 1141–1160 of 2520Page 58 of 100
NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1141

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If M is the mean of n observations x1 - k, x2 - k, x3 - k, _ _ _, xn - k, where k is any real number, then what is the mean of x1, x2, x3, _ _ _, xn?

  1. AOption (a)

    M

  2. BOption (b)

    M + k

  3. COption (c)

    M - k

  4. DOption (d)

    kM

View answer & explanation

Correct answer

Answer: (b)B.

M + k

Why this answer is correct

Concept:

For the observations \(\rm x_1 ,x_2 ,x_3 ,..........,x_n\)

Mean (\(\rm\overline x\)) = \(\rm {x_1 +x_2 + x_3 +..........+x_n\over n}\)

Calculation:

Given mean of the n observation is M

M = \(\rm x_1 - k +x_2 - k+ x_3 - k +..........+x_n-k\over n\)

M = \(\rm (x_1 +x_2 + x_3 +..........+x_n)-nk\over n\)

M = \(\rm {x_1 +x_2 + x_3 +..........+x_n\over n}-{nk\over n}\)

M = \(\rm {x_1 +x_2 + x_3 +..........+x_n\over n}\) - k

M + k = \(\rm {x_1 +x_2 + x_3 +..........+x_n\over n}\)

∴ The required mean of x1, x2, x3, _ _ _, xn = M + k

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1142

National Defence Academy (NDA) 2021• Mathematics• Mathematics

What is the sum of deviations of the variate values 73, 85, 92, 105, 120 from their mean?

  1. AOption (a)

    -2

  2. BOption (b)

    -1

  3. COption (c)

    0

  4. DOption (d)

    5

View answer & explanation

Correct answer

Answer: (c)C.

0

Why this answer is correct

Concept:

For the observations \(\rm x_1 ,x_2 ,x_3 ,..........,x_n\)

Mean (\(\rm\overline x\)) = \(\rm {x_1 +x_2 + x_3 +..........+x_n\over n}\)

Deviation of the observation from the mean = xi - \(\rm\overline x\)

where i ∈ [1, n]

Calculation:

Mean = \(\rm {73+85+92+105+120\over 5}\)

\(\rm\overline x\) = \(\rm {475\over 5}\) = 95

ValueMeanDeviation of the value
739573 - 95 = -22
859585 - 95 = -10
929592 - 95 = -3
10595105 - 95 = 10
12095120 - 95 = 25

 

The sum of the deviation = -22 +(-10) + (-3) + 10 + 25 = 0

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1143

National Defence Academy (NDA) 2021• Mathematics• Mathematics

Let x be the HM and y be the GM of two positive numbers m and n. If 5x = 4y, then which one of the following is correct?

  1. AOption (a)

    5m = 4n

  2. BOption (b)

    2m = n

  3. COption (c)

    4m = 5n

  4. DOption (d)

    m = 4n

View answer & explanation

Correct answer

Answer: (d)D.

m = 4n

Why this answer is correct

Concept:

For any n numbers \(\rm x_1 ,x_2 ,x_3 ,..........,x_n\)

GM = \(\rm \sqrt[n]{x_1\times x_2\times x_3\times ..........\times x_n}\)

HM = \(\rm {\left[{1\over x_1}+{1\over x_2}+{1\over x_3}+..........+{1\over x_n}\over n\right]}^{-1}\)

Calculation:

GM of m and n = y

\(\rm \sqrt{mn}\) = y

HM of m and n = x

\(\rm {\left[{1\over m}+{1\over n}\over 2\right]}^{-1}\) = x

\(\rm {2mn\over m+ n}\) = x

Now given 5x = 4y

\(\rm {10mn\over m+ n}\) = \(\rm 4\sqrt{mn}\)

\(\rm 10\sqrt{mn}\) = 4(m + n)

100mn = 16(m2 + n2 + 2mn)

16m2 + 16n2 - 68mn = 0

\(\rm16{m\over n}+16{n\over m} -68= 0\)

Let \(\rm m\over n\) = t

16t + \(\rm 16\over t\) - 68 = 0

16t2 - 68t + 16 = 0

t = \(\rm {68 \pm \sqrt{(-68)^2-4(16)(16)} \over 2\times16}\)

t = \(\rm {68 \pm \sqrt{4624-1024} \over 32}\)

t = \(\rm {68 \pm60\over 32}\) = 4 or \(\rm 1\over 4\)

\(\rm m\over n\) = 4 or \(\rm 1\over 4\)

∴ m = 4n or n = 4m

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1144

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If the mean of a frequency distribution is 100 and the coefficient of variation is 45%, then what is the value of the variance?

  1. AOption (a)

    2025

  2. BOption (b)

    450

  3. COption (c)

    45

  4. DOption (d)

    4.5

View answer & explanation

Correct answer

Answer: (a)A.

2025

Why this answer is correct

Concept:

Coefficient of variation = \(\rm\text{Standard Deviation} \over\text{ Mean}\)

Variance = (Standard Deviation)2

Calculation:

Given coefficient of variation = 45% = 0.45

And mean = 100

As Coefficient of variation = \(\rm\text{Standard Deviation} \over\text{ Mean}\)

0.45 = \(\rm\text{Standard Deviation} \over100\)

Standard Deviation = 100 × 0.45

SD = 45

∴ Variance = 452 = 2025

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1145

National Defence Academy (NDA) 2021• Mathematics• Mathematics

Let two events A and B be such that P(A) = L and P(B) = M. Which one of the following is correct?

  1. AOption (a)

    \(P(A|B)<\dfrac{L+M-1}{M}\)

  2. BOption (b)

    \(P(A|B)>\dfrac{L+M-1}{M}\)

  3. COption (c)

    \(P(A|B)\ge\dfrac{L+M-1}{M}\)

  4. DOption (d)

    \(P(A|B)=\dfrac{L+M-1}{M}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(P(A|B)\ge\dfrac{L+M-1}{M}\)

Why this answer is correct

Concept:

\(\rm P(A|B)=\dfrac{P(A∩ B)}{P(B)}\)

P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

Calculation:

Given P(A) = L and P(B) = M

Here P(A ∪ B) ≤ 1 (∵ Max value of probability is 1)

P(A) + P(B) - P(A ∩ B) ≤ 1

L + M - P(A ∩ B) ≤ 1

P(A ∩ B) ≥ L + M - 1

Now \(\rm P(A|B)=\dfrac{P(A∩ B)}{P(B)}\)

\(\rm P(A|B)\geq\dfrac{L+M-1}{M}\)

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1146

National Defence Academy (NDA) 2021• Mathematics• Mathematics

For which of the following sets of numbers do the mean, median and mode have the same value?

  1. AOption (a)

    12, 12, 12, 12, 24

  2. BOption (b)

    6, 18, 18, 18, 30

  3. COption (c)

    6, 6, 12, 30, 36

  4. DOption (d)

    6, 6, 6, 12, 30

View answer & explanation

Correct answer

Answer: (b)B.

6, 18, 18, 18, 30

Why this answer is correct

Calculation:

Option 2 has 

Mean = \(6+18+18+18+30\over5 \) = 18

Median = 18 {As it is the middle value of the arrangement}

Mode = 18 {As it appears the most}

∴ Mean = Median = Mode

Option 2 is correct

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1147

National Defence Academy (NDA) 2021• Mathematics• Mathematics

The mean of 12 observations is 75. If two observation are discarded, then the mean of the remaining observations is 65. What is the mean of the discarded observations?

  1. AOption (a)

    250

  2. BOption (b)

    125

  3. COption (c)

    120

  4. DOption (d)

    Cannot be determined due to insufficient data

View answer & explanation

Correct answer

Answer: (b)B.

125

Why this answer is correct

Calculation:

Let the discarded values be x and y 

Sum of all the 12 values = 12 × mean

S1 = 12 × 75 = 900

Sum of the 10 values after discarding = 10 × new mean

S2 = 10 × 65 = 650

Now 650 + x + y = 900

x + y = 900 - 650 = 250

Average of x and y =\(\rm x+y\over2\)

Average = \(\rm 250\over2\) = 125

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1148

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If k is one of the roots of the equation x(x + 1) + 1 = 0, then what is its other root?

  1. AOption (a)

    1

  2. BOption (b)

    -k

  3. COption (c)

    k2

  4. DOption (d)

    -k2

View answer & explanation

Correct answer

Answer: (c)C.

k2

Why this answer is correct

Concept:

For a quadratic equation ax2 + bx + c = 0

The sum of the roots = \(\rm-{b\over a}\)

The product of the roots = \(\rm{c\over a}\)

Calculation:

Let the other root be β  

Given equation is x(x + 1) + 1 = 0

⇒ x2 + x + 1 = 0

a = 1, b = 1 and c = 1

As k is the root of the equation

⇒ k2 + k + 1 = 0

⇒ k2 = -1 - k     .....(i)

The sum of the roots = \(-{1\over1}\) = -1

⇒ β + k = -1

⇒ β = -1 - k      .....(ii)

From equation (i) and (ii), we get

⇒ β = k2

∴ The other root = k2

 

<sup>

 </sup>

Given equation is x(x + 1) + 1 = 0

Factor of (x2 + x + 1) = 0

\({\rm{x}} = {\rm{;}}\frac{{ - 1{\rm{;}} \pm {\rm{;}}\sqrt {{1^2} - 4{\rm{;}} \times 1{\rm{;}} \times 1} }}{{2{\rm{;}} \times 1}} = {\rm{;}}\frac{{ - 1{\rm{;}} \pm {\rm{i}}\sqrt 3 }}{2}\)

\(⇒ {\rm{x}} = {\rm{;}}\frac{{ - 1 + {\rm{i}}\sqrt 3 }}{2}{\rm{;or;;}}\frac{{ - 1 - {\rm{i}}\sqrt 3 }}{2}\)

⇒ x = ω or ω2

Consider k = ω 

∴ The other root = ω2 = k2

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1149

National Defence Academy (NDA) 2021• Mathematics• Mathematics

The geometric mean of a set of observations is computed as 10. The geometric mean obtained when each observations xi is replaced by \(\rm 3x_i^4\) is

  1. AOption (a)

    810

  2. BOption (b)

    900

  3. COption (c)

    30000

  4. DOption (d)

    81000

View answer & explanation

Correct answer

Answer: (c)C.

30000

Why this answer is correct

Concept:

If GM of series xi = a, then 

  • GM of λxi = λa
  • GM of \(\rm {x_i}^n\) = an

Calculation:

Given GM of xi = 10

Then GM of \(\rm {x_i}^4\) = 104 = 10000

Now GM of \(\rm 3{x_i}^4\) = 3 × 10000 = 30000

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1150

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If \(\rm P(A\cup B)=\dfrac{5}{6}, P(A\cap B)=\dfrac{1}{3}:and:P(\bar A)=\dfrac{1}{2}\), then which of the following is/are correct?

  1. A and B are independent events.
  2. A and B are mutually exclusive events.

Select the correct answer using the code given below.

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (a)A.

1 only

Why this answer is correct

Concept:

For any events A and B

  • P(A ∪ B) = P(A) + P(B) - P(A ∩ B)
  • Events are independent if P(A ∩ B) = P(A) × P(B)
  • Events are mutually exclusive if P(A ∩ B) = 0
  • P(\(\rm\bar A\)) = 1 - P(A)

 

Calculation:

Given \(\rm P(A\cup B)=\dfrac{5}{6}, P(A∩ B)=\dfrac{1}{3}:and:P(\bar A)=\dfrac{1}{2}\)

P(A) = 1 - \(1\over2\) = \(1\over2\)

P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

\(\rm{5\over6} = {1\over2}+P(B)-{1\over3}\)

P(B) = \(2\over3\)

Given P(A ∩ B) ≠ 0

P(A ∩ B) = P(A) × P(B) = \({1\over2}\times{2\over3} = {1\over3}\)

∴ A and B are independent events and are not mutually exclusive events.

Only statement 1 is correct.

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1151

National Defence Academy (NDA) 2021• Mathematics• Mathematics

The average of a set of 15 observations is recorded, but later it is found that for one observation, the digit in the tens place was wrongly recorded as 8 instead of 3. After correcting the observation, the average is

  1. AOption (a)

    reduced by \(\dfrac{1}{3}\)

  2. BOption (b)

    increased by \(\dfrac{10}{3}\)

  3. COption (c)

    reduced by \(\dfrac{10}{3}\)

  4. DOption (d)

    reduced by 50

View answer & explanation

Correct answer

Answer: (c)C.

reduced by \(\dfrac{10}{3}\)

Why this answer is correct

Formula Used:

Average = \(\frac{X_i}{n}\)

Here, Xi is the sum of observation and n is the number of reading

Calculation:

Let the average initially be x

Then the sum of the observations = 15x

As the tens digit is recorded 5 more, therefore the sum is 50 more than the correct sum. 

Correct sum of the observations = 15x - 50

The correct average = \(\rm 15x-50\over15\) = x - \(10\over3\)

∴ The average is reduced by \(10\over3\)

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1152

National Defence Academy (NDA) 2021• Mathematics• Mathematics

A coin is tossed twice. If E and F denote occurrence of head on first toss and second toss respectively, then what is P(E ∪ F) equal to?

  1. AOption (a)

    \(\dfrac{1}{4}\)

  2. BOption (b)

    \(\dfrac{1}{2}\)

  3. COption (c)

    \(\dfrac{3}{4}\)

  4. DOption (d)

    \(\dfrac{1}{3}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\dfrac{3}{4}\)

Why this answer is correct

Concept:

For the independent events A and B, 

  • P(A ∪ B) = P(A) + P(B) - P(A ∩ B)
  • P(A ∩ B) = P(A) × P(B)

 

Calculation:

Probability of getting head on first toss P(E) = \(1\over2\)

Probability of getting head on first toss P(F) = \(1\over2\)

Since the events are independent P(E ∩ F) = P(E) × P(F)

P(E ∩ F) = \(1\over2\) × \(1\over2\) = \(1\over4\)

P(E ∪ F) = P(E) + P(F) - P(E ∩ F)

P(E ∪ F) = \(1\over2\) + \(1\over2\) - \(1\over4\)

P(E ∪ F) = \(3\over4\)

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1153

National Defence Academy (NDA) 2021• Mathematics• Mathematics

In a binomial distribution, the mean is \(\dfrac{2}{3}\) and variance is \(\dfrac{5}{9}\). What is the probability that random variable D = 2?

  1. AOption (a)

    \(\dfrac{5}{36}\)

  2. BOption (b)

    \(\dfrac{25}{36}\)

  3. COption (c)

    \(\dfrac{25}{54}\)

  4. DOption (d)

    \(\dfrac{25}{216}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\dfrac{25}{216}\)

Why this answer is correct

Concept:

For the binomial distribution of probability

  • Mean = np
  • Variance = npq

Where n is the total number of cases, p is probability of favorable cases and q is probability of unfavorable cases(1 - p)

Calculation:

Given mean = np = \(2\over3\) and variance = npq = \(5\over9\)

\(\rm variance \over mean\) = \({5\over9}\over{2\over3}\)

q = \(5\over6\)

p = 1 - q = \(1\over6\)

np = \(2\over3\)

n = 4

∴ The probability that random variable D = 2

P = \(\rm {^nC_2}p^2q^{n-2}\)

P = \(\rm {^4C_2}\times\left({1\over6}\right)^2\times\left({5\over6}\right)^2\)

P = \(\rm 6\times{1\over36}\times{25\over36}\)

P = \(25\over216\)

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1154

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If the mode of the scores 10, 12, 13, 15, 15, 13, 12, 10, x is 15, then what is the value of x?

  1. AOption (a)

    10

  2. BOption (b)

    12

  3. COption (c)

    13

  4. DOption (d)

    15

View answer & explanation

Correct answer

Answer: (d)D.

15

Why this answer is correct

Concept:

The mode of the n observation is the number that has the highest frequency.

Calculation:

Given scores 10, 12, 13, 15, 15, 13, 12, 10, x and mode = 15

Frequency of score '12' = 2

Frequency of score '15' = 2

As Mode = 15, its frequency has to be more than 2, if the value of x = 15

Then the frequency of score 15 = 3 & is highest.

So, x should be '15'

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1155

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If A and B are two events such that \(\rm P(A)=\dfrac{3}{4} :and: P(B)=\dfrac{5}{8}\), then consider the following

Statements:

  1. The minimum value of P(A ∪ B) is \(\dfrac{3}{4}.\)
  2. The maximum value of P(A ∩ B) is \(\dfrac{5}{8}.\)

Which of the above statements is/are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (c)C.

Both 1 and 2

Why this answer is correct

Concept:

For any two events A and B

  • The minimum value of P(A ∪ B) = max [P(A), P(B)]
  • The maximum value of P(A ∩ B) = min [P(A), P(B)]

 

Calculation:

Given \(\rm P(A)=\dfrac{3}{4} :and: P(B)=\dfrac{5}{8}\)

P(A) > P(B)

∴ The minimum value of P(A ∪ B) = \(\dfrac{3}{4}\)

The maximum value of P(A ∩ B) = \(\dfrac{5}{8}\)

Hence both 1 and 2 are correct

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1156

National Defence Academy (NDA) 2021• Mathematics• Mathematics

What is the derivative of ex with respect to xe?

  1. AOption (a)

    \(\dfrac{xe^x}{ex^e}\)

  2. BOption (b)

    \(\dfrac{e^x}{x^e}\)

  3. COption (c)

    \(\dfrac{xe^x}{x^e}\)

  4. DOption (d)

    \(\dfrac{e^x}{ex^e}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\dfrac{xe^x}{ex^e}\)

Why this answer is correct

Concept:

The differentiation of f(x) wrt g(x) = \(\rm f'(x)\over g'(x)\)

f(x) and g(x) be the function of x

Calculation:

f(x) = ex 

f'(x) = \(\rm {d\over dx}e^x = e^x\)

g(x) = xe 

g'(x) = \(\rm {d\over dx}x^e = ex^{e-1}\)

Required derivative D = \(\rm f'(x)\over g'(x)\)

D = \(\rm e^x\over ex^{e-1}\) = \(\rm xe^x\over ex^e\)

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1157

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If a differentiable function f(x) satisfies \(\mathop {\lim }\limits_{x \to - 1} \dfrac{f(x)+1}{x^2-1}=-\dfrac{3}{2}\) then what is \(\mathop {\lim }\limits_{x \to - 1} f(x)\) equal to?

  1. AOption (a)

    \(-\dfrac{3}{2}\)

  2. BOption (b)

    -1

  3. COption (c)

    0

  4. DOption (d)

    1

View answer & explanation

Correct answer

Answer: (b)B.

-1

Why this answer is correct

Calculation:

Given: \(\rm\mathop {\lim }\limits_{x \to - 1} \dfrac{f(x)+1}{x^2-1}=-\dfrac{3}{2}\)

Denominator becomes zero on putting the limits

∴ The numerator must also be zero to have a certain value of the given limit, i.e.,

\(\rm \mathop {\lim }\limits_{x \to - 1} [f(x)+1]=0\)

\(\rm \mathop {\lim }\limits_{x \to - 1} f(x)\) + 1 = 0

\(\rm \mathop {\lim }\limits_{x \to - 1} f(x)\) = -1

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1158

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If the function \(\rm f\left( x \right) = \left{ {\begin{array}{*{20}{c}} {a + bx,;;}&{x < 1}\ {5,}&{x = 1}\ {b - ax,}&{x > 1} \end{array}} \right.\) is continuous, then what is the value of (a + b)?

  1. AOption (a)

    5

  2. BOption (b)

    10

  3. COption (c)

    15

  4. DOption (d)

    20

View answer & explanation

Correct answer

Answer: (a)A.

5

Why this answer is correct

Concept:

For the function to be continuous:

LHL = RHL = f(x)

Where LHL = \(\rm\mathop {\lim }\limits_{α \to 0}\)f(x - α) and RHL = \(\rm\mathop {\lim }\limits_{α \to 0}\)f(x + α)

Calculation:

Given that f(x) is continuous function

LHL = f(x) = RHL

\(\rm\mathop {\lim }\limits_{α \to 0}\) f(1 - α) = f(1)

\(\rm\mathop {\lim }\limits_{α \to 0}\) [a + b(1 - α)] = 5

\(\rm\mathop {\lim }\limits_{α \to 0}\) [a + b - bα] = 5

a + b = 5

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1159

National Defence Academy (NDA) 2021• Mathematics• Mathematics

Consider the following statements in respect of the function f(x) = sin x:

  1. f(x) increases in the interval (0, π).
  2. f(x) decreases in the interval \(\left(\dfrac{5\pi}{2},3\pi\right).\)

Which of the above statements is/are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (b)B.

2 only

Why this answer is correct

Concept:

The function f(x):

  • Increases for the values x of having f'(x) > 0
  • Decreases for the values of x having f'(x) < 0

 

Calculation:

Given function is f(x) = sin x

∴ f'(x) = cos x

For x ∈ (0, \(π\over2\)), cos x > 0 and for x ∈ (\(π\over2\), π), cos x < 0

∴ For x ∈ (0, π), f(x) increases to max at \(π\over2\) and then decreases to min at π.

For x ∈ \(\left(\dfrac{5\pi}{2},3\pi\right)\), cos x < 0

∴ f(x) decreases for x ∈ \(\left(\dfrac{5\pi}{2},3\pi\right)\)

statement 2 is correct Only.

NDA Mathematics 18 Apr 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1160

National Defence Academy (NDA) 2021• Mathematics• Mathematics

What is the domain of the function f(x) = 3x?

  1. AOption (a)

    (-∞, ∞)

  2. BOption (b)

    (0, ∞)

  3. COption (c)

    [0, ∞)

  4. DOption (d)

    (-∞, ∞) - {0}

View answer & explanation

Correct answer

Answer: (a)A.

(-∞, ∞)

Why this answer is correct

Concept:

The domain is the set of all possible value of x which have a finite value of f(x).

Calculation:

Given function f(x) = 3x

The function will have a finite value for all x ∈ (-∞, ∞)

Mistake PointsThe range of the given function will be from (0,∞). 0, when x = -∞, and ∞ when x = ∞.

How should I prepare for Quantitative Aptitude in National Defence Academy (NDA)?

Solving previous year questions by topic helps identify recurring concepts, question patterns, and frequently asked sub-topics. Attempt each question, review the step-by-step explanation, and note down key formulas or facts for quick revision before the exam.